# AI-enhanced tuning of quantum dot Hamiltonians toward Majorana modes

### v1

@@ I Introduction
Majorana zero modes (MZMs) are emergent quasiparticles predicted to obey non-Abelian statistics, enabling topologically protected quantum information processing [Alicea2012, Beenakker2013, Sau2010, Oreg2010]. Their realization in hybrid superconductor-semiconductor systems requires precise control over local parameters such as chemical potentials, inter-dot couplings, and spin-orbit interactions [Leijnse2012, Lutchyn2018]. Small deviations from the so-called sweet-spot conditions of the Kitaev chain (KC) [Leumer2021] can destroy the topological gap and delocalize MZMs [Kitaev2001]. Moreover, it is not trivial to distinguish between MZMs and different nontopological zero-bias peaks  [Aguado2021, Sasaki2000, Pikulin2012, Kells2012]. Consequently, achieving robust Majorana states experimentally remains a major challenge, particularly in the presence of fabrication disorder and parameter noise.
To address this challenge, Fulga et al. [Fulga2013] proposed an adaptive tuning protocol for superconductor-proximitized chain of quantum dots (QDs) that simulates a KC generalized on non-uniform parameters distribution across dots, trying to demonstrate how gate-controlled voltages and superconducting phases can be adjusted to reach a topologically nontrivial regime. In this paper we revisit this problem trying to automate it using deep neural estimators. More recently, further works indicate that through appropriate tuning, it is possible to realize a KC with emerged MZMs using QDs coupled through elastic co-tunneling and crossed Andreev reflection  [Leijnse2012, Sau2012, Dvir2023, Bordin2025]. All these proposals utilize coupling with more readily available $s$-wave superconductors; however, there are also proposals employing $p$-wave superconductors [Mohseni2023, Ezawa2024].
At the same time, a different strategy for machine learning (ML)-assisted autotuning of QD-based quantum simulators [Borsoi2024, Mills2019, Shandilya2025] using transport measurements [Darulov2020, Lunczer2019, Zwolak2023, roux2025] is gaining significant interest. The use of transport measurements in the form of conductance maps [Donarini2024, Shandilya2025] to derive insights about the Hamiltonian system [Blonder1982, grepkowa2023, Wang2017, Gebhart2023, Thamm2024, Taylor2024disorderlearning, Taylor2025analysis] appears to be a promising path towards the automatic tuning of Hamiltonian parameters. In particular, inverting the measured conductance matrices to determine electrostatic potential disorder, using evolutionary optimization [Thamm2024], supervised neural networks (NNs) [Taylor2024disorderlearning, vandriel2024, Taylor2025analysis], or both [Taylor2025disordermitigation], is a natural first step that can allow precise tuning of Hamiltonian parameters towards MZMs. Although these methods already allow for successful mitigation of disorder in nanowires [Taylor2025disordermitigation, Thamm2024], they rely on indirect, non-differentiable cost functions and heuristic evolutionary searches, rather than learning the broader behavior of the underlying physical system.
@@
In contrast, here we introduce an unsupervised physics-informed [Raissi2019, kovachki2023neural] NN-based auto-tuning (PINNAT) model with Majorana physics embedded in the loss function, and show that by proper training the model can memorize various structures on conductance maps which enables to build an efficient quantum simulator autotuning system. Our AI-enhanced adaptive tuning framework directly extends QDs simulator by leveraging vision transformer (ViT) [dosovitskiy2021] NN architecture. The scheme of the proposed approach is presented in Fig. 1 with information flow during the model training (black arrows) and evaluation (inference, blue arrows).
@@ II Model
Following a lattice model from Ref. [Fulga2013], the effective superconductor-proximitized Rashba-Zeeman Hamiltonian for a chain of spinful single-level QDs can be expressed as follows
@@
$+\Bigl.\left(t_{n}(e^{i\bm{\lambda}_{n}\cdot\bm{\sigma}})_{ss^{\prime}}c_{n,s}^{\dagger}c_{n+1,s^{\prime}}+\mathrm{h.c.}\right)\Bigr],$ (1)
with the on-site potentials $\mu_{n}$, inter-dot hopping amplitudes $t_{n}$ modulated by (Rashba) spin-orbit vector $\bm{\lambda_{n}}=\lambda_{n}(\sin\rho_{n}\cos\xi_{n},\sin\rho_{n}\sin\xi_{n},\cos\rho_{n})$, Zeeman energy $V_{\mathrm{Z}}$, proximity-induced superconducting ($s$-wave) pairing $\Delta_{n}e^{i\phi_{n}}$, and $n$ describing dot number while $s$ being spin degree of freedom. The values of those parameters are then carefully adjusted to have MZMs emerge in the system. One set of tuned (yet uniform) parameters (which we call reference parameters, $P_{0}$) for $N=3$ QDs chain includes: $\mu=0.6\,\mathrm{meV}$, $t=0.25\,\mathrm{meV}$, $\lambda=0.27\,\pi$, $\rho=\xi=\frac{\pi}{2}$, $V_{\mathrm{Z}}=0.5\,\mathrm{meV}$, and $\Delta=0.25\,\mathrm{meV}$. The Rashba length $\lambda=0.27\,\pi$ was tuned so that at $\mu=0.6$ meV two energy levels touch at zero energy (c.f. Fig. 5(a)). Also note that some parameters, i.e., $\{\mu_{n}$, $t_{n}$, $\lambda_{n}\}$ (7 in total) are easier to tune – electrically (via local gating) than the others: the global Zeeman field $V_{\mathrm{Z}}$ and the proximity-induced superconducting gap $\{\Delta_{n}\}$.
To train the ViT model in physics-informed manner, we introduce a differentiable quasi-metric $\mathcal{M}$, quantifying how close a given Hamiltonian is to MZM regime. The proposed phenomenology combines several physical indicators: edge-state localization, zero-energy spectral weight, and parity (electron-hole) symmetry. For a detailed $\mathcal{M}$ definition, see the Methods section. We also note that we tested the popular Majorana polarization [Sticlet2012, Maska2017, Szumniak2017] measure, directly related to the parity operator, but it yielded unsatisfactory training results. This is because it fails to discriminate topologically trivial zero modes that localize in the central QD or anti-localize in the left and right QDs (see the Supplementary Information of Ref. [Bordin2025] for examples of such states). We define our measure in a way that also discriminates these states.
@@
for the Hamiltonian with tuned parameters $H(P^{\prime})=H(P\,+\,\delta P)$. Additionally, to force the PINNAT to predict smallest possible correction $\delta P$ to the parameters $P$, an extra regularization term $\alpha\langle\delta P\rangle^{2}$ is included, with a factor $\alpha=0.1$. Training details, including parameters sampling used to generate synthetic training set of conductance maps, are described in Methods section.
@@ III Results
To present the PINNAT performance in parameter tuning, we plot the values of $\mathcal{M}$, disturbing selected three-QDs Hamiltonian parameters. Specifically, we shift (detune) the value of the given parameters (the rest is kept default, $P_{0}$) – obtaining $P$ – and verify what corrections $\delta P$ are predicted by the NN models. Then the $\mathcal{M}(H(P+\delta P))$ value is plotted. In Fig. 2(b) and (c) we show how $\mathcal{M}$ changes with uniform (i.e., the same for each QD) parameter shift for models tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ and $\{\mu_{n},V_{\mathrm{Z}}\}$, respectively. The values for $\mathcal{M}(H(P))$, i.e. before correction, are presented as reference in column (a). For example, in $t$ vs $\mu$ plot we shift parameters to $t_{1}=t_{2}=t$ and $\mu_{1}=\mu_{2}=\mu_{3}=\mu$ obtaining $P$ (column (a)), then predict correction $\delta P$ and present $\mathcal{M}(H(P+\delta P))$ in respective positions in columns (b) and (c). Additionally, in column Fig. 2(a), orange dashed curves mark the analytical conditions: (1) $\Delta=\sqrt{V^{2}_{\mathrm{Z}}-\mu^{2}}$ for the existence of two zero modes in decoupled QDs, and (2) $\lambda=\arctan(\frac{\mu}{\Delta})$ corresponding to the KC sweet spot (see the Supplementary Notes (SN) for details).
@@
Interestingly, the model that adjusts $\{\mu_{n},V_{\mathrm{Z}}\}$ – Fig. 2(c) is slightly more effective in proposing corrections, both in terms of parameter range that can be effectively corrected and the probability of MZMs emergence, as indicated by higher values of $\mathcal{M}$. This behavior is specifically worth highlighting as the model does not modify the $t_{n}$, $\lambda_{n}$ and $\Delta$ parameters, while still possessing the ability to effectively amend them by updating $\mu_{n}$ and $V_{\mathrm{Z}}$. These results show that $V_{\mathrm{Z}}$ plays a more important role in controlling the system than the hoppings $t_{n}$ and SOI $\lambda_{n}$. This is consistent with observations known from long chains [stanescu2024, Leumer2021], where the magnetic field is essential to compete with the superconducting pairing to drive the system into the topological phase. However, it should be noted that when the hopping parameters $t$, $\lambda$ are close to zero in the first and second rows of Fig. 2(c) the model likewise selects symmetric zero modes localized at the edge QDs with a high $\mathcal{M}$; nevertheless in this case (in the limit $t\rightarrow 0$) the system is clearly in a trivial phase [stanescu2024].
@@
Figure 4: Iterative autotuning procedure for (a) model adjusting $\{\mu_{n},t_{n},\lambda_{n}\}$, and (b) adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$. In the first (left) plot, we present $\mathcal{M}$ map before tuning. Then, in the center, parameters are tuned with a single step of NN corrections. Finally, in the last (right) plot, there is a map plotted for parameters tuned within 10 subsequent steps of NN corrections.
@@
Figure 5: Three-QD chain Hamiltonian $H$ as a function of the offset $\mu$ added to local potentials: $\mu_{n}\rightarrow\mu_{n}+\mu-0.6$ meV (dashed vertical line marks the reference $\mu=0.6\,\mathrm{meV}$). Additionally $\mathcal{M}$ metric is plotted, and eigenvalues are colored (left column) with edge occupations – 1 means that the state is localized on the edges, while 0 means that state is localized in the center dot, or (right column) electron-hole symmetry – calculated as difference between density of electrons and holes. Plots are presented for (a) reference parameters; (b) modified $\lambda_{1}$ and $\lambda_{2}$; (c) parameters from (b) but with NN-tuned $\{\mu_{n},t_{n},\lambda_{n}\}$; (d) parameters from (b) with NN-tuned $\{\mu_{n},V_{\mathrm{Z}}\}$.
@@
Adding a noise to $\lambda_{n}$ (by setting $\lambda_{0}=0.62\,\pi$ and $\lambda_{1}=0.41\,\pi$) – presented in Fig. 5(b) – results in reduced $\mathcal{M}$ value and vanishing MZMs. Then the noisy parameter can be corrected using either of the two presented NN models. Tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ – presented in Fig. 5(c) – makes MZMs reappear with desired properties, i.e., they are localized on the edges, electron-hole symmetry is preserved, and the edge states are slightly gapped from the rest of the spectrum. On the other hand, tuning $\{\mu_{n},V_{\mathrm{Z}}\}$ – Fig. 5(d) – not only leads to restoring MZMs at the exact point in the tuned parameter space with the larger topological gap but also increases the probability of observing MZMs in a wider range of $\mu$ values. An additional example of analogous autotuning, but for the $t$ parameter is provided in SN.
@@ IV Discussion
These findings align with and extend current efforts in ML-assisted control of Majorana platforms [Thamm2024, Taylor2025disordermitigation, vandriel2024]. Most notably, the recent work of Taylor and Das Sarma, Ref. [Taylor2025disordermitigation] introduced a ViT-based framework for mitigating strong disorder in Majorana nanowires. Their study demonstrated that disorder can be compensated – sometimes restoring topology in systems that originally show no nontrivial regions – using conductance-based NN inference combined with evolutionary optimizer (CMA-ES). Our approach differs in several important respects. First, instead of relying on a second-stage optimizer such as CMA-ES, our network learns Hamiltonian behaviors manifested by transport measurements to directly predict parameter corrections, making the tuning process faster and conceptually closer to a experimental tuning protocol. Second, whereas Ref. [Taylor2025disordermitigation] focuses primarily on restoring scattering-invariant and LDOS-based indicators in bulk nanowires, our work addresses QD chains, a complementary Majorana platform with discrete QDs and gate-defined tunnel couplings between them. Third, our method integrates NN physics-informed via differentiable quasi-metric $\mathcal{M}$ capturing edge localization, electron-hole symmetry, and spectral proximity to zero energy; this contrasts with supervised learning to pretict topology-based indicators.
@@
Importantly, the proposed PINNAT scheme is fully physics-informed: MZM metric $\mathcal{M}$, together with the the Hamiltonian family definition and eigensolver implementation, is explicitly encoded in the structure of the NN loss function. This approach aligns with the recent strong trend toward designing NNs that respect the underlying physics of the problem [Bronstein2021], for example in studies of quantum entanglement [Pawlowski2024, Krawczyk2024], quantum control [Norambuena2024] or in the optimization of quantum tomography protocols [Krawczyk2025].
While our current proof-of-concept study is restricted mainly to a three-QDs system, we show the ability to scale the proposed method to 7-QDs chain in the SN. With some minor adjustments in the training procedure we claim that it is possible to have the PINNAT model trained to propose significant corrections in the larger systems, maintaining its ability to autotune the Hamiltonian parameters beyond the training regime.
@@
V Conclusions
VI Methods
Proposed method uses conductance maps as high-dimensional visual inputs for the ViT model. The conductance $G$ is calculated using the $S$-matrix formalism in the wide-band limit [Donarini2024] via the Weidenmüller formula [Bordin2025, christiansen2009] for $N=3$ QDs chain:

### v3

@@ I Introduction
Majorana zero modes (MZMs) are emergent quasiparticles predicted to obey non-Abelian statistics, enabling topologically protected quantum information processing [1, 2, 41, 34]. Their realization in hybrid superconductor-semiconductor systems requires precise control over local parameters such as chemical potentials, inter-dot couplings, and spin-orbit interactions [22, 29]. Small deviations from the so-called sweet-spot conditions of the Kitaev chain (KC) [24] can destroy the topological gap and delocalize MZMs [17]. Moreover, it is not trivial to distinguish between MZMs and different nontopological zero-bias peaks  [51, 40, 36, 16]. Consequently, achieving robust Majorana states experimentally remains a major challenge, particularly in the presence of fabrication disorder and parameter noise.
To address this challenge, Fulga et al. [14] proposed an adaptive tuning protocol for superconductor-proximitized chain of quantum dots (QDs) that simulates a KC generalized on non-uniform parameters distribution across dots, trying to demonstrate how gate-controlled voltages and superconducting phases can be adjusted to reach a topologically nontrivial regime. In this paper we revisit this problem trying to automate it using deep neural estimators. More recently, further works indicate that through appropriate tuning, it is possible to realize a KC with emerged MZMs using QDs coupled through elastic co-tunneling and crossed Andreev reflection [23, 22, 42, 12, 5]. All these proposals utilize coupling with more readily available $s$-wave superconductors; however, there are also proposals employing $p$-wave superconductors [32, 13].
At the same time, a different strategy for machine learning (ML)-assisted autotuning of QD-based quantum simulators [6, 31, 43] using transport measurements is gaining significant interest [9, 28, 55, 39, 27]. The use of transport measurements in the form of conductance maps [10, 43] to derive insights about the system Hamiltonian [3, 18, 53, 15, 50, 49, 48] appears to be a promising path towards the automatic tuning of its parameters. In particular, inverting the measured conductance matrices to determine electrostatic potential disorder, using evolutionary optimization [50], supervised neural networks (NNs) [49, 52, 48], or both [47], is a natural first step that can allow precise tuning of Hamiltonian parameters towards MZMs. Although these methods already allow for successful mitigation of disorder in nanowires [47, 50], they rely on indirect, non-differentiable cost functions and heuristic evolutionary searches, rather than learning the broader behavior of the underlying physical system.
@@
In contrast, here we introduce an unsupervised physics-informed [38, 19] NN-based auto-tuning (PINNAT) model with Majorana physics embedded in the loss function, and show that by proper training the model can memorize various structures on conductance maps which enables to build an efficient quantum simulator autotuning system. Our AI-enhanced adaptive tuning framework directly extends QDs simulator by leveraging vision transformer (ViT) [11] NN architecture. The scheme of the proposed approach is presented in Fig. 1 with information flow during the model training (black arrows) and evaluation (inference, blue arrows).
@@ II Model
Following a lattice model from Ref. [14], the effective superconductor-proximitized Rashba-Zeeman Hamiltonian for a chain of spinful single-level QDs can be expressed as follows
@@
$+\Bigl.\left(t_{n}(e^{i\boldsymbol{\lambda}_{n}\cdot\boldsymbol{\sigma}})_{ss^{\prime}}c_{n,s}^{\dagger}c_{n+1,s^{\prime}}+\mathrm{h.c.}\right)\Bigr],$ (1)
with the on-site potentials $\mu_{n}$, inter-dot hopping amplitudes $t_{n}$ modulated by (Rashba) spin-orbit vector $\boldsymbol{\lambda_{n}}=\lambda_{n}(\sin\rho_{n}\cos\xi_{n},\sin\rho_{n}\sin\xi_{n},\cos\rho_{n})$, Zeeman energy $V_{\mathrm{Z}}$, proximity-induced superconducting ($s$-wave) pairing $\Delta_{n}e^{i\phi_{n}}$, and $n$ describing dot number while $s$ being spin degree of freedom. The values of those parameters are then carefully adjusted to have MZMs emerge in the system. One set of tuned (yet uniform) parameters (which we call reference parameters, $P_{0}$) for $N=3$ QDs chain includes: $\mu=0.6\,\mathrm{meV}$, $t=0.25\,\mathrm{meV}$, $\lambda=0.27\,\pi$, $\rho=\xi=\frac{\pi}{2}$, $V_{\mathrm{Z}}=0.5\,\mathrm{meV}$, and $\Delta=0.25\,\mathrm{meV}$. The Rashba length $\lambda=0.27\,\pi$ was tuned so that at $\mu=0.6$ meV two energy levels touch at zero energy (c.f. Fig. 5(a)). Also note that some parameters, i.e., $\{\mu_{n}$, $t_{n}$, $\lambda_{n}\}$ (7 in total) are easier to tune – electrically (via local gating) than the others: the global Zeeman field $V_{\mathrm{Z}}$ and the proximity-induced superconducting gap $\{\Delta_{n}\}$.
To train the ViT model in physics-informed manner, we introduce a differentiable quasi-metric $\mathcal{M}$, quantifying how close a given Hamiltonian is to MZM regime. The proposed phenomenology combines several physical indicators: edge-state localization, zero-energy spectral weight, and parity (electron-hole) symmetry. For a detailed $\mathcal{M}$ definition, see the Methods section. We also note that we tested the popular Majorana polarization [45, 30, 46] measure, directly related to the parity operator, but it yielded unsatisfactory training results. This is because it fails to discriminate topologically trivial zero modes that localize in the central QD or anti-localize in the left and right QDs (see the Supplementary Information of Ref. [5] for examples of such states). We define our measure in a way that also discriminates these states.
@@
for the Hamiltonian with tuned parameters $H(P^{\prime})=H(P\,+\,\delta P)$. Additionally, to force the PINNAT to predict smallest possible correction $\delta P$ to the parameters $P$, an extra regularization term $\alpha\langle\delta P\rangle^{2}$ is included, with a factor $\alpha=0.1$. Training details, including parameters sampling used to generate synthetic training set of conductance maps, are described in the Methods section.
@@ III Results
To present the PINNAT performance in parameter tuning, we plot the values of $\mathcal{M}$, disturbing selected three-QDs Hamiltonian parameters. Specifically, we shift (detune) the value of the given parameters (the rest is kept default, $P_{0}$) – obtaining $P$ – and verify what corrections $\delta P$ are predicted by the NN models. Then the $\mathcal{M}(H(P+\delta P))$ value is plotted. In Fig. 2(b) and (c) we show how $\mathcal{M}$ changes with uniform (i.e., the same for each QD) parameter shift for models tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ and $\{\mu_{n},V_{\mathrm{Z}}\}$, respectively. The values for $\mathcal{M}(H(P))$, i.e. before correction, are presented as reference in column (a). For example, in $t$ vs $\mu$ plot we shift parameters to $t_{1}=t_{2}=t$ and $\mu_{\mathrm{L}}=\mu_{\mathrm{C}}=\mu_{\mathrm{R}}=\mu$ obtaining $P$ (column (a)), then predict correction $\delta P$ and present $\mathcal{M}(H(P+\delta P))$ in respective positions in columns (b) and (c). Additionally, in column Fig. 2(a), orange dashed curves mark the analytical conditions: (1) $\Delta=\sqrt{V^{2}_{\mathrm{Z}}-\mu^{2}}$ for the existence of two zero modes in decoupled QDs, and (2) $\lambda=\arctan(\frac{\mu}{\Delta})$ corresponding to the KC sweet-spot (see the Appendix A for details).
@@
Interestingly, the model that adjusts $\{\mu_{n},V_{\mathrm{Z}}\}$, presented in Fig. 2(c), is slightly more effective in proposing corrections, both in terms of parameter range that can be effectively corrected and the probability of MZMs emergence, as indicated by higher values of $\mathcal{M}$. This behavior is specifically worth highlighting as the model does not modify the $t_{n}$, $\lambda_{n}$ and $\Delta$ parameters, while still possessing the ability to effectively amend them by updating $\mu_{n}$ and $V_{\mathrm{Z}}$. These results show that $V_{\mathrm{Z}}$ plays a more important role in controlling the system than the hoppings $t_{n}$ and SOI $\lambda_{n}$. This is consistent with observations known from Rashba wires [1, 44]. Having a (perpendicular) magnetic field removes the so-called fermion doubling problem by opening a gap (at $k=0$). The size of this gap increases with the Zeeman field, which means that for larger Zeeman splitting there is a wider range of $\mu_{n}$ for which the nontrivial phase is present. This is what we observe. However, it should be noted that when the hopping parameters $t$, $\lambda$ are close to zero in the first and second rows of Fig. 2(c) the model likewise selects symmetric zero modes localized at the edge QDs with a high $\mathcal{M}$; nevertheless in this case (in the limit $t\rightarrow 0$) the system is clearly in a trivial phase [44]. For $\lambda=0$ there is no possibility of realizing $p$-wave pairing in the KC, and the entire model breaks down.
@@
One way to mitigate this behavior is to include regularization terms in the training phase that push these parameters (t and $\lambda$) away from zero. However, to simplify the training procedure, we decided not to include it in this work.
@@
Figure 4: Iterative autotuning procedure for (a) model adjusting $\{\mu_{n},t_{n},\lambda_{n}\}$, and (b) adjusting $\{\mu_{n},V_{\mathrm{Z}}\}$. In the first (left) plot, we present $\mathcal{M}$ map before tuning. Then, in the center, parameters are tuned with a single step of $N$ corrections. Finally, in the last (right) plot, there is a map plotted for parameters tuned within 10 subsequent steps of $N$ corrections.
@@
Figure 5: Three-QD chain Hamiltonian $H$ as a function of the offset $\mu$ added to local potentials: $\mu_{n}\rightarrow\mu_{n}+\mu-0.6$ meV (dashed vertical line marks the reference $\mu=0.6\,\mathrm{meV}$). Additionally, the $\mathcal{M}$ metric is plotted, and eigenvalues are colored (left column) with edge occupations $O_{edge}^{(i)}$, or (right column) electron-hole symmetry $S_{\mathrm{eh}}^{(i)}$ defined in the Methods section. Plots are presented for (a) reference parameters; (b) modified $\lambda_{1}$ and $\lambda_{2}$; (c) parameters from (b) but with NN-tuned $\{\mu_{n},t_{n},\lambda_{n}\}$; (d) parameters from (b) with NN-tuned $\{\mu_{n},V_{\mathrm{Z}}\}$. Orange arrow points to the trivial zero modes.
@@
Adding a noise to $\lambda_{n}$ (by setting $\lambda_{1}=0.62\,\pi$ and $\lambda_{2}=0.41\,\pi$) – presented in Fig. 5(b) – results in reduced $\mathcal{M}$ value and vanishing MZMs. Then the noisy parameter can be corrected using either of the two presented NN models. Tuning $\{\mu_{n},t_{n},\lambda_{n}\}$ – presented in Fig. 5(c) – makes MZMs reappear with desired properties, i.e., they are localized on the edges, electron-hole symmetry is preserved, and the edge states are slightly gapped from the rest of the spectrum. On the other hand, tuning $\{\mu_{n},V_{\mathrm{Z}}\}$ – Fig. 5(d) – not only leads to restoring MZMs at the exact point in the tuned parameter space with the larger topological gap but also increases the probability of observing MZMs in a wider range of $\mu$ values. An additional example of analogous autotuning, but for the $t$ parameter is provided in the Appendix B.
Importantly, Fig. 5 shows that the $\mathcal{M}$ metrics achieve high values only for the topological states. Specifically, trivial zero modes like the Andreev bound state (ABS) [26, 37] indicated by the orange arrow in Fig. 5(a) and Fig. 5(b) do not increase the $\mathcal{M}$ value. Since the PINNAT model is trained to increase $\mathcal{M}$, only non-trivial zero-energy states emerge after parameter tuning, highlighting the robustness of the proposed approach in distinguishing between trivial and topological states.
@@ IV Discussion
These findings align with and extend current efforts in ML-assisted control of Majorana platforms [50, 47, 52]. Most notably, the recent work of Taylor and Das Sarma, Ref. [47] introduced a ViT-based framework for mitigating strong disorder in Majorana nanowires. Their study demonstrated that disorder can be compensated – sometimes restoring topology in systems that originally show no nontrivial regions – using conductance-based NN inference combined with evolutionary optimizer (CMA-ES). Our approach differs in several important respects. First, instead of relying on a second-stage optimizer such as CMA-ES, our network learns Hamiltonian behaviors manifested by transport measurements to directly predict parameter corrections, making the tuning process faster and conceptually closer to a experimental tuning protocol. Second, whereas Ref. [47] focuses primarily on restoring scattering-invariant and LDOS-based indicators in bulk nanowires, our work addresses QD chains, a complementary Majorana platform with discrete QDs and gate-defined tunnel couplings between them. Third, our method integrates NN physics-informed via differentiable quasi-metric $\mathcal{M}$ capturing edge localization, electron-hole symmetry, and spectral proximity to zero energy; this contrasts with supervised learning to predict topology-based indicators.
It is instructive to compare the parameter regimes identified by PINNAT with theoretical sweet-spot relations discussed in Appendix A. The parameter sets obtained in our NN-based autotuning follow the same qualitative trends, namely Zeeman energy exceeding the inter-dot couplings and remaining comparable to $\sqrt{\Delta^{2}+\mu_{n}^{2}}$, while spin-orbit coupling following the expected sweet-spot relation with onsite potential. These relations are satisfied at the level of order-of-magnitude and averaged parameters, rather than exactly on each site, reflecting the fact that PINNAT optimizes the metric $\mathcal{M}$ instead of enforcing analytical constraints explicitly. A detailed comparison is provided in Appendix D.
@@
Importantly, the proposed PINNAT scheme is fully physics-informed: MZM metric $\mathcal{M}$, together with the Hamiltonian family definition and eigensolver implementation, is explicitly encoded in the structure of the NN loss function. This approach aligns with the recent strong trend toward designing NNs that respect the underlying physics of the problem [7], for example in studies of quantum entanglement [35, 21], quantum control [33], or in the optimization of quantum tomography protocols [20]. Moreover, we argue that similar approaches based on conductance map inputs could be developed to optimize a broad class of physics-informed metrics. In particular, since direct estimation of Hamiltonian parameters from conductance maps appears feasible [Pawłowski2026], a wide range of parameter-tuning problems could potentially be addressed using PINNAT-like models. Tuning the QD system into the single-electron (Coulomb blockade) regime could constitute a standard application in this context.
While our current proof-of-concept study is restricted mainly to a three-QDs system, we show the ability to scale the proposed method to a 7-QDs chain in the Appendix C. With some optimization adjustments in the training procedure, we claim that it is possible to have the PINNAT model trained to propose significant corrections in the larger systems, maintaining its ability to autotune the Hamiltonian parameters beyond the training regime.
Although our demonstrations focus on QD chains, the PINNAT framework is not restricted to this platform. Since proximitized semiconductor nanowires are routinely modeled using discretized lattice Hamiltonians of the same type [29, 24], the proposed approach could be extended to nanowire systems—albeit with a limited number of lattice sites, as discussed in Appendix C—by replacing the underlying Hamiltonian and adapting transport formalism.
@@
V Methods
Proposed method uses conductance maps as high-dimensional visual inputs for the ViT model. The conductance $G$ is calculated using the $S$-matrix formalism in the wide-band limit [10] via the Weidenmüller formula [5, 8] for $N=3$ QDs chain: